OpenStudy (anonymous):

I hope someone can help me with this, because I have no idea how to do it! and if you could please show me how to do it would be amazing! Thanks! A missile is launched from the ground. Its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the missile is 110 feet in the air; after 2 seconds, it is 200 feet in the air. Find the height, in feet, of the missile after 4 seconds in the air.

OpenStudy (anonymous):

let h(x)=a x^2 + bx +c

OpenStudy (anonymous):

we need to find a,b,c from the given conditions

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

when x=0 h =0 means c=0 when x=1 h=110 means 110 = a+b when x=2 h=200 means 200 = 4a+2b so we need to solve a+b= 110 and 4a+2b =200

OpenStudy (anonymous):

i hope u can do further

OpenStudy (anonymous):

im not sure how to do these, but I'll try it

OpenStudy (anonymous):

i got 320

OpenStudy (anonymous):

Thank you matricked!

OpenStudy (anonymous):

a+b =110 2a+b= 100 subtracting the aboe two eq we have -a= 10 or a=-10 hence b= 120

OpenStudy (anonymous):

we need h when x=4 so h= -10*16 + 120*4=480-160 =320